### v2

# Can Generalist Vision Language Models (VLMs) Rival Specialist Medical VLMs? Benchmarking and Strategic Insights

Abstract
Background: Vision–Language Models (VLMs) have shown promise in automating image diagnosis and interpretation in clinical settings. However, developing specialist medical VLMs requires substantial computational resources and carefully curated datasets, and it remains unclear under which conditions generalist and specialist medical VLMs each perform best.
Methods: This paper introduces , the first unified benchmark and evaluation framework designed for direct, fair comparisons of paired generalist and specialist medical VLMs from the same model families. Our benchmark is comprehensive, assessing 18 models across radiology, pathology, dermatology, and ophthalmology on 144 diagnostic and 72 visual-question answering (VQA) tasks. focusing on assessing both in-domain (ID) and out-of-domain (OOD) performance, with off-the-shelf and parameter-efficient fine-tuning (e.g., linear probing, LoRA). Diagnostic classification tasks were evaluated using AUROC, while VQA tasks were assessed with BLEU-1, ROUGE-L, Exact Match, F1 Score, and GPT-based semantic scoring, covering both open- and closed-ended formats.
Results: As expected, off-the-shelf specialist medical VLMs generally outperformed generalist VLMs on ID tasks given their pretraining. However, with lightweight fine-tuning, generalist VLMs achieved superior performance in most of ID task evaluations and demonstrated better generalization on OOD tasks in the majority of comparisons. Fine-tuning required only updating small amounts of parameters associated with full medical pretraining. In contrast, fine-tuned specialist medical VLMs do not show comparable performance even with the same level of fine-tuning. is fully open-sourced at https://github.com/ubc-tea/MedVLMBench.
Conclusions: This study highlights the complementary strengths of specialist medical and generalist VLMs. Specialists remain valuable in modality-aligned use cases, but we find that efficiently fine-tuned generalist VLMs can achieve comparable or even superior performance in most tasks, particularly when transferring to unseen or rare OOD medical modalities. These results suggest that generalist VLMs, rather than being constrained by their lack of specialist medical pretraining, may offer a scalable and cost-effective pathway for advancing clinical AI development.
Can Generalist Vision Language Models (VLMs) Rival Specialist Medical VLMs?
Benchmarking and Strategic Insights
1 Introduction
The rapid evolution of artificial intelligence (AI) in healthcare has accelerated the development of vision–language models (VLMs) for medical imaging—large-scale systems trained to align visual and textual modalities [lu2025integrating]. These models promise a generalist capability across tasks ranging from clinical image-text processing to diagnostic image interpretation by leveraging extensive unsupervised pretraining over large and diverse datasets. Recent advances have produced domain-specialist medical VLMs 111We define specialist medical VLMs as the VLMs pretrained or continue-pretrained primarily on medical corpora. , such as BioMedCLIP [zhang2023biomedclip], LLaVA-Med [li2023llava], and Med-Gemma [sellergren2025medgemma], adopting extensive pretraining on multimodal medical corpora spanning radiology, pathology, ophthalmology, and dermatology. Despite their promise, such specialist pipelines demand substantial computational resources and curated data at scale, rendering them expensive and technically challenging to develop [bommasani2021opportunities].
In many real-world clinical applications, AI models are expected to excel at specific tasks, particularly within the data distributions in which they are deployed. The reliance on broad generalist models may not always align with the needs of high-stakes, domain-specific applications in medical imaging. While the prevailing assumption is that larger models trained on diverse medical datasets yield superior generalization, this raises another hypothesis that a domain-agnostic VLM model, with minimal fine-tuning, could rival or even outperform an expensive specialist medical model on certain medical imaging tasks. Furthermore, claims of strong adaptability in specialist medical VLMs are questionable in out-of-domain as limited evaluation is conducted on their unseen medical imaging modalities
In contrast, generalist VLMs-such as CLIP [radford2021learning], LLaVA [liu2023visual], and Gemma [team2025gemma]-though not originally developed for medicine, are widely available, pretrained on massive general-domain data, and demonstrate strong transferable priors. Crucially, these models can often be adapted for medical applications through lightweight fine-tuning (e.g., linear probing, parameter-efficient methods) at a fraction of the cost. This contrast raises a central research question: Are large-scale, specialist medical VLMs truly necessary, or can generalist VLMs—when efficiently adapted—achieve comparable performance on critical medical imaging tasks?
To systematically tackle the central problem above by strategically comparing the generalist and specialist VLMs, this paper formulates the following three research questions (RQs):
However, neither existing medical VLM benchmarks nor related studies are designed to address all of these questions. They are limited in three critical ways [xu2025lingshu, royer2024multimedeval, jeong2024medical]: 1) Overlooking Adaptation Effects: These studies mainly focus on evaluating pre-trained models, overlooking the fact that in real-world applications these models are often adapted (e.g., fine-tuned) for specific downstream tasks, and ignoring the substantial performance changes that can occur after this adaptation. 2) Inadequate Comparison Framework: They focus on evaluating different medical models against each other, rather than conducting a systematic comparison between specialist VLMs and their generalist counterparts in the same model family 222Model variants can be grouped into families based on their backbone, as shown in Appendix A. 3) Lack of task diversity: The current evaluation lacks a robust, comprehensive testing protocol with a consistent in-domain (ID) vs. out-of-domain (OOD) framework. This makes it challenging to measure how different VLMs generalize across varied tasks and conditions.
To provide plausible answers to the above research questions, this paper introduces , a benchmark that can fill the current research gaps. Firstly, is designed to directly assess VLMs in their off-the-shelf (OTS) form and after lightweight adaptation (e.g., linear probing and parameter-efficient fine-tuning). Secondly, further couples performance evaluation with a simple cost proxy to allow fair, like-for-like comparisons by comparing generalist VLMs with their specialist medical counterparts in the same model family and globally. Specifically, encompasses two main VLM paradigms: contrastive models and generative multimodal large language models. Each paradigm includes both generalist (e.g., CLIP [radford2021learning], LLaVA [liu2023visual], Gemma [team2025gemma]) and specialist medical representatives (e.g., BioMedCLIP [zhang2023biomedclip], LLaVA-Med [li2023llava], Med-Gemma [sellergren2025medgemma]). And it focuses on two representative categories of tasks in medical imaging: disease diagnosis and visual question answering (VQA). These tasks capture both discriminative and generative aspects of VLMs. Finally, to reflect the dual needs of clinical deployment: excelling on ID scenarios, where training and test modalities align, and generalizing to OOD settings, where models encounter unseen modalities or tasks, we extensively evaluate on both settings. Furthermore, systematically evaluates models across ten diverse medical datasets (six diagnostic, four VQA). The protocol is structured to directly probe the core of each question by comparing OTS versus fine-tuned VLMs, and ID versus OOD generalization, encompassing 144 disease diagnostic and 72 VQA scenarios (Figure 2).
Our evaluation protocol yields direct answers to these questions. For RQ1, we find that OTS specialist medical VLMs predictably outperform generalist VLMs on their respective ID tasks. However, addressing RQ2, we demonstrate this advantage is fragile: lightweight fine-tuning (e.g., linear probing) and parameter-efficient fine-tuning (PEFT) on medical data allows generalist VLMs to not only match but frequently surpass the performance of these specialist models. Critically, for RQ3, we find that this fine-tuning process on OOD datasets also confers superior generalization, as fine-tuned generalist VLMs prove more robust and effective on new tasks than their specialist medical counterparts, which can exhibit performance degradation and over-specialization.
In summary, contributes in the following three aspects:
1.
It provides the first systematic benchmarking of representative generalist and specialist medical VLMs across ten datasets, covering 144 diagnostic and 72 VQA scenarios.
2.
It directly addresses three fundamental research questions, offering new insights into the relative strengths of generalist versus specialist models.
3.
It further investigates how lightweight adaptation can effectively help the generalist model adapt specialist tasks, revealing a practical and cost-efficient path for practical clinical VLM deployment.
2 Method
The overall pipeline of is shown in Figure 1. is structured to capture both ID and OOD performance across two major task families: diagnostic classification and VQA. Within this setup, we examine models in both their OTS form and after PEFT, enabling a controlled assessment of adaptability and robustness. The following subsections describe the datasets used, the models evaluated, and the evaluation protocols adopted in the paper.
2.1 Datasets
utilizes 10 large medical datasets across diverse imaging modalities (radiology, pathology, dermatology, ophthalmology) and clinical tasks for robust evaluation (Fig. 1a). Disease diagnosis datasets include CheXpert (CXP) [irvin2019chexpert], Camelyon [bandi2018detection], HAM10000 (HAM) [tschandl2018ham10000], PAPILA [papineni2002bleu], GF3300 [luo2024harvard] , and FairVLMed [luo2024fairclip]. Medical VQA datasets included VQA-RAD [lau2018dataset], PathVQA [he2020pathvqa], SLAKE [liu2021slake], and FairVLMed [luo2024fairclip], each containing both closed-form (yes/no) and open-form question–answer pairs. FairVLMed is simultaneously used for both tasks. GPT-4 is used to design one closed-form question and one open-form question from each clinical report in FairVLMed for VQA evaluation. Across all datasets, no data points were missing, and we relied exclusively on the existing data. Dataset statistics are reported in Appendix B (Tables 3 and 4) and details of the VQA dataset construction are provided in the Appendix C.
ID and OOD datasets. Considering that modern large specialist medical VLMs are typically trained on one or more modalities across a broad range of organs and tasks, we categorize downstream tasks and evaluations as ID or OOD at the modality level. For each model, datasets are designated as ID if their primary modality was represented during pretraining, and OOD otherwise. The composition and proportions of training data for most generalist models are not clearly disclosed. We considered that, although the large, web-scale pretraining data for generalist VLMs may include some medical content-incidentally or intentionally-these models are not explicitly trained for or specialized in the medical domain. Notably, most evaluated medical models are further pre-trained on the backbone of their corresponding generalist model and therefore inherit the same medical priors, if any. Therefore, this is unlikely to confound our within-family comparisons in OOD analyses. Moreover, the evaluation data for the downstream tasks we tested is not reported as part of their pretraining. Consequently, we treat these medical tasks as OOD for the generalist models.
2.2 Models
Models were grouped into two major categories based on learning paradigm: contrastive and generative VLMs. Within each category, both the generalist and the corresponding specialist medical counterpart were included. The choice of models was guided by their prevalence in the literature, diversity of pre-training sources, and relevance to current medical AI applications.
Contrastive VLMs. Contrastive VLMs map images and text into a common embedding space using contrastive learning to enable zero-shot diagnosis. Our evaluation is structured around two core architectural families. First, the CLIP family is examined, comparing the foundational web-trained CLIP [radford2021learning] and BLIP2 [li2023blip] against its numerous medical-domain counterparts, which include models specialized for radiology (MedCLIP [wang2022medclip], PubMedCLIP [eslami2023pubmedclip]), pathology (PLIP [huang2023visual]), and broad biomedical data (BioMedCLIP [zhang2023biomedclip]). Second, assess the SigLIP family, represented by the generalist SigLIP [zhai2023sigmoid] and its medical adaptation, MedSigLIP [sellergren2025medgemma]. By pairing generalist and specialist models within each family, we can directly investigate the impact of domain-specific contrastive pre-training strategies.
Generative VLMs. Generative VLMs integrate image and text through multimodal transformers capable of producing free-form, human-like responses. For our analysis, we selected models from several prominent families, each represented by a generalist version and its specialist medical counterparts. These include the LLaVA family, comprising the generalist LLaVA-1.5 [liu2023visual] and the medical pre-trained LLaVA-Med [li2023llava]; the Gemma family, including the generalist Gemma-3 [team2025gemma] and specialist MedGemma [sellergren2025medgemma]; and the Qwen family, including the Qwen-VL [bai2025qwen2] and the specialist counterpart Lingshu [xu2025lingshu]. To benchmark these against the broader state-of-the-art, we also include the flagship close-weight generalist models from SenseTime (InternVL3 [zhu2025internvl3]), OpenAI (o3), and Google (Gemini 2.5 Pro) in the OTS settings.
Together, this cohort of models covers a wide spectrum of architectural designs and pre-training methodologies, with detailed specifications provided in Appendix A.
2.3 Metrics
For disease diagnosis tasks, the area under the receiver operating characteristic curve (AUROC) was the primary metric. For VQA tasks, we follow standard practices [sharma2021survey, xu2025lingshu, li2023llava, nath2025vila] and evaluate overall closed- and open-ended questions using token-level matching metrics that quantify overlap between model responses and reference answers, including exact match, accuracy, precision, recall, and F1 score. Notebly, we also follow [liu2023visual] and employ the GPT score as a language model–based measure of semantic answer quality. For open-ended questions, we further report BLEU-1 (unigram precision) [papineni2002bleu] and ROUGE-L (longest common subsequence recall) [lin2004rouge].
Because OTS models often diverge from references in response structure and style, token-matching metrics can be biased. Therefore, we report analyses based on the GPT score in the main text, which are comparatively style-invariant and focus on semantic correctness. Implementation details for the GPT score and the full set of metric results are provided in the Appendix C.
2.4 Evaluation and statistical methods
Evaluation settings. Two evaluation modes were considered. In the OTS setting, contrastive VLMs were applied directly in a zero-shot manner using their native similarity scoring between image and text embeddings, while generative VLMs were prompted with task-specific instructions and their outputs were scored against references. In the PEFT setting, different adaptation strategies were applied depending on the model type. For contrastive VLMs, two approaches were used: (i) linear probing (LP), in which a shallow linear classifier is trained on top of frozen image–text embeddings; and (ii) LoRA, which introduces trainable low-rank adapters within the encoder blocks while keeping the remaining parameters fixed. In the main paper, we focus on LP given its lightweight suitability for clinical deployment, and relegate the full set of results to the Appendix D. For generative VLMs, fine-tuning involved both the bridge module (typically the multilayer perceptron that maps vision encoder features into the language space) and the language decoder itself, with LoRA applied to the decoder to enable efficient adaptation. This protocol ensures a consistent comparison between generalist and specialist medical VLMs across architectures and domains, while keeping the number of trainable parameters to be light-weighted.
Research questions. To address the three guiding research questions (RQ1–RQ3), performance differences were summarized as “RQ gaps” below:
RQ1 gap: Difference between the generalist and the specialist medical model when both are used in the OTS setting on ID datasets.
RQ2 gap: Difference between the fine-tuned generalist model and the OTS specialist model on ID datasets.
RQ3 gap: Difference between the fine-tuned generalist model and the fine-tuned specialist medical model when both are evaluated on OOD datasets.
Statistical methods. To estimate uncertainty and support robust comparisons, nonparametric bootstrap resampling with 1,000 iterations was applied to both diagnostic and VQA tasks. Samples were drawn with replacement to calculate all metrics in Sec. 2.3. Percentile-based 95% confidence intervals and standard deviations were derived from the empirical distributions. This design provides assumption-light evaluation of generalist and specialist VLMs under both ID and OOD conditions.
3 Results
Figure 2 summarizes the numerical benchmark outcomes across all tasks in a heatmap, which our analysis below consistently refers to. Overall, the specialists’ medical VLMs (green) consistently outperform generalist VLMs (blue) in OTS usage.
We further conduct systematic comparison analysis on the results in Figure 3 and Figure 4 to address our research questions. Subsection 3.1 analyzes the family-wise gaps presented in Figure 3. Next, subsection 3.2 performs a global comparative analysis on individual datasets to compare the adapted generalist VLMs with the best-performing specialist medical VLMs, which is presented in Figure 4. Both subsections are structured according to our three RQs 1–3 to quantify the performance difference between generalist and specialist medical VLMs across diagnosis and VQA tasks.
Together, these two views allow us to address all our research questions. The complete raw data underlying these figures are included in Appendix D, with Tables 5–7 reporting diagnostic results and Tables 8–27 reporting VQA results.
3.1 Comparison between generalist VLMs and their counterpart specialist medical VLMs
Figure 3 visualizes the RQ gaps between the best-performing generalist and their counterpart specialist medical VLMs within each family, which support systematic comparisons under consistent architectural and modeling conditions. A negative RQ1 gap (green line) reflects a specialist advantage in the OTS setting; a positive RQ2 gap (blue line) indicates that generalists surpass specialists after light-weight adaptation; and a positive RQ3 gap (pink line) denotes superior generalization of generalists on OOD tasks.
RQ1 Setting - Specialist VLMs dominate in OTS, ID settings.
The RQ1 gaps (green) are consistently negative, showing that specialists medical hold a strong advantage without adaptation, which reflects the benefit of modality-specific pretraining. For instance, Figure 3a shows that in radiology (CXP) and pathology (Camelyon), specialist medical CLIP-families set the performance ceiling. Concrete results from Figure 2 illustrate this: MedCLIP achieves 90.60% (CI: 84.78%, 95.39%) AUROC on CXP, far above CLIP, which achieves 40.95% (CI: 31.68%, 50.67%). Likewise, MedGemma outperforms Gemma3 on VQA-RAD by nearly 12 points in GPT Score. These examples confirm that specialist priors directly translate into higher OTS accuracy on ID datasets.
RQ2 Setting - Lightweight fine-tuning flips the advantage toward generalists.
The RQ2 gaps (blue) in Figure 3a–b turn positive almost uniformly (except certain VLMs on CXP), revealing that generalist VLMs rapidly close and often surpass the specialist lead after adaptation. For example, BLIP2 overtakes PLIP on Camelyon once tuned, jumping from a near-random OTS baseline to above 98% AUROC (Figure 2). Similarly, in VQA, LLaVA-1.5 surpasses LLaVA-Med on PathVQA after tuning. These cases demonstrate that broad visual–textual pretraining in generalists can be efficiently specialized, potentially obviating the need for expensive domain-specific pretraining to achieve state-of-the-art performance.
RQ3 Setting - Generalists generalize more effectively to OOD tasks.
The RQ3 gaps (pink) highlight generalists’ superior adaptability. Figure 3a–b shows that tuned generalists outperform specialists on almost all OOD datasets. For instance, in dermatology (HAM), BLIP2 and SigLIP exceed 87% AUROC, while specialist medical VLM like MedCLIP remain below 80% (Figure 2). In multimodal VQA (SLAKE), Qwen2.5-VL reaches 86.3% (CI: 84.43%, 88.15%) GPT Score, well above tuned specialist medical counter parts plateauing at 72–74%. Moreover, one can observe that certain specialist medical VLMs (e.g., Lingshu) sometimes exhibit degraded performance after adaptation. Similar studies suggest that this also stems from their poor generalization, as overly trained specialist models is found to struggle to adapt when adapted under even the same settings . These observations align with recent literature [springer2025overtrained]. These findings suggest that specialists, while strong in their pretraining domain, overfit to modality-specific features, whereas generalists retain flexible representations that transfer across new modalities.
3.2 Comparison between generalist VLMs and the best-performing specialist medical VLMs
Building on the controlled within-family analysis in Sec. 3.1, Figure 4 broadens the scope to a global comparison that benchmarks fine-tuned generalist VLMs against best specialist models, regardless of model family. This global evaluation offers additional evidence for the effectiveness of fine-tuning generalist models.
The results indicate that the key insights from Sec. 3.1 persist in this broader comparative analysis. For RQ1, top-2 specialized medical VLMs consistently stand out in OTS ID scenarios as shown by green dashed lines. For RQ2, nine of the twelve fine-tuned generalist diagnostic baselines (comprising three bars each across the CXP, Camelyon, HAM, and PAILA datasets) surpass the best OTS specialized medical models, with all twelve outperforming the second-ranked specialist models. Similarly, ten of the fifteen VQA baselines (spanning VQA-RAD, PathVQA, and SLAKE) exceed the performance of the best OTS specialized medical VLM. For RQ3, fine-tuned generalists consistently meet or exceed the performance of best fine-tuned specialized medical models in OOD contexts across diverse domains (radiology, pathology, dermatology, and ophthalmology) and tasks (diagnosis and VQA). In diagnosis (Figure 4a), eight out of eighteen tuned generalist baselines exceed even the best-tuned specialist medical OOD models, while eleven outperform the second-best. In VQA, with the exception of InternVL, all generalist VLMs surpass the strongest medical OOD VLMs (FairVLMed).
4 Conclusion
benchmarks and shows that generalist and specialist medical VLMs have distinct strengths in medical imaging. When the clinical task closely matches the model’s pretraining modality-for example, radiology-focused models on chest X-ray interpretation or pathology-tuned models on histopathology datasets-specialist medical VLMs exploit domain-specific priors to detect modality-dependent patterns that generalist models may overlook, thus achieving strong performance. By contrast, generalist VLMs demonstrate superior adaptability in OOD and cross-modality contexts. When adapted for OOD downstream tasks, they consistently outperform specialist medical VLMs across a majority of diagnostic and VQA datasets. This advantage likely stems from their broad pretraining on diverse visual-linguistic corpora, which enables more flexible abstraction and knowledge transfer to new medical tasks.
The strong OTS performance of close-weights generalist VLMs like OpenAI’s o3 and Google’s Gemini 2.5 Pro further demonstrates this superior generalizability, as shown in Figure 2. These models have already demonstrate proficiency on clinical tasks [arora2025healthbench]. However, we do not extend the discussion to these closed-weight models because they cannot be included in the fine-tuning experiments for RQ2 and RQ3. Furthermore, they were trained on extremely broad, heterogeneous datasets that likely contain extensive downstream medical modalities, but the recipe of medical corpus not explicitly mentioned in their technical reports, making controlled expriment difficult.
is not without limitations. As with any benchmark, has natural scope boundaries. Our study evaluates a broad and finite set of models and tasks, and the results reflect the specific models and datasets currently included. Such scope of our evaluation enhances clarity and reproducibility and does not diminish the value of as a rigorous, systematic benchmark for comparing generalist and specialist medical VLMs in controlled, diverse settings. Importantly, is designed to be readily extensible: additional models, modalities, and medical datasets can be incorporated with minimal effort. To ensure fair and robust comparisons as the field evolves, we recommend applying the comparison protocols when extending to unevaluated modalities, emerging datasets, or future generations of VLM architectures, rather than directly extrapolating the present findings.
Appendix A Vision-Language Models
| VLM | Domain | Training modality | Training dataset(s) | Public | Model size |
| CLIP family | |||||
| CLIP[radford2021learning] | Generalist | Image–Text (web) | 400 M WebImageText pairs | ✓ | 151.28M |
| BLIP-2[li2023blip] | Generalist | Image–Text (web) | COCO[lin2014microsoft], Visual Genome[krishna2017visual], CC3M[sharma2018conceptual], CC12M[changpinyo2021conceptual], SBU[ordonez2011im2text], LAION-400M[schuhmann2021laion] (129 M pairs) | ✓ | 1172.03M |
| MedCLIP[wang2022medclip] | Specialist Medical | Chest X-ray pairs | CheXpert[irvin2019chexpert], MIMIC-CXR[johnson2019mimic] (570 k pairs) | ✓ | 136.62M |
| PubMedCLIP[eslami2023pubmedclip] | Specialist Medical | Radiology images | ROCO (80 k images from PubMed Central) | ✓ | 151.28M |
| PLIP[huang2023visual] | Specialist Medical | Pathology images | OpenPath (21 k Twitter pathology pairs) | ✓ | 151.28M |
| BioMedCLIP[zhang2023biomedclip] | Specialist Medical | Multi-modality medical figures | PMC-15M (15.3 M figure–caption pairs) | ✓ | 195.90M |
| SigLIP family | |||||
| SigLIP[zhai2023sigmoid] | Generalist | Image–Text (web) | WebLI[chen2022pali] (10 B pairs) | ✓ | 203.16M |
| MedSigLIP[sellergren2025medgemma] | Specialist Medical | Mixed (general + medical) | SigLIP and Gemma-3 mixtures; Med-Gemini corpus | ✓ | 878.30M |
| VLM | Domain | Vision encoder | LLM backbone | Training dataset(s) | Public | Model size |
| LLaVA family | ||||||
| LLaVA-1.5 [liu2023visual] | Generalist | CLIP ViT-L/336 | Vicuna 7B | MLP projector; 1.2 M visual instructions (GPT-4, academic VQA) | ✓ | 7.38B |
| LLaVA-Med [li2023llava] | Specialist Medical | CLIP ViT-L/336 | Mistral-7B | Two-stage curriculum on 15 M PMC figure–caption pairs + GPT-4 self-instruct | ✓ | 7.90B |
| Qwen-VL family | ||||||
| Qwen-2-VL [wang2024qwen2] | Generalist | Dynamic-res ViT | Qwen-2 LLM | Web-scale image-text pairs, videos; SFT with dialogue, grounding, and agent data | ✓ | 8.29B |
| Qwen-2.5-VL [bai2025qwen2] | Generalist | Dynamic-res ViT | Qwen-2.5 LLM | Web-scale data; Multimodal Pre-Training, Long-Context Pre-Training, SFT, tool-use alignment | ✓ | 8.29B |
| Lingshu [xu2025lingshu] | Specialist Medical | Qwen-VL ViT | Qwen-2.5 (base) | Shallow align deep align med-instruction tune Reinforcement learning with verifiable rewards | ✓ | 8.29B |
| InternVL family | ||||||
| InternVL 3 [zhu2025internvl3] | Generalist | InternViT-L (V2PE) | Qwen2.5 7B | Native joint multimodal pre-training SFT + mixed-preference opt. | ✓ | 7.94B |
| Gemma family | ||||||
| Gemma 3 [team2025gemma] | Generalist | SigLIP ViT-B (400 M, frozen) | Gemma-3 4B | Distillation on multi-modal+text mix; vision tokens | ✓ | 4.30B |
| MedGemma [sellergren2025medgemma] | Specialist Medical | MedSigLIP ViT | Gemma-3 4B | Multi-stage med corpus; frozen encoder, 256 vision tokens | ✓ | 4.97B |
| OpenAI family | ||||||
| o3 [openai2025o3o4mini] | Generalist | Proprietary multimodal encoder | OpenAI (o3) | Proprietary multimodal pre-training; RL-driven alignment; instruction tuning | ✓ | NA |
| Gemini family | ||||||
| 2.5 Pro [comanici2025gemini] | Generalist | Proprietary multimodal encoder | Gemini 2.5 Pro | Joint multimodal pre-training; instruction tuning and preference optimization | ✓ | NA |
Appendix B Datasets
| Dataset | Modality | Train | Test | Download URL |
| SLAKE | CT, MRI, and X-Ray | 5980 | 1061 | https://huggingface.co/datasets/BoKelvin/SLAKE |
| PathVQA | Pathology | 26373 | 6719 | https://huggingface.co/datasets/flaviagiammarino/path-vqa |
| VQARAD | Radiology | 2,244 | 451 | https://huggingface.co/datasets/flaviagiammarino/vqa-rad |
| FairVLMed | SLO fundus | 17994 | 3994 | https://ophai.hms.harvard.edu/datasets/harvard-fairvlmed10k |
| Dataset | Modality | Train | Test | Download URL |
| CheXpert | Radiology | 223414 | 234 | https://stanfordmlgroup.github.io/competitions/chexpert/ |
| GF3300 | Fundus photography | 2706 | 594 | https://ophai.hms.harvard.edu/datasets/harvard-gf3300/ |
| PAPILA | Color retinal fundus | 364 | 56 | https://www.nature.com/articles/s41597-022-01388-1 |
| HAM10000 | Dermatology | 8137 | 1811 | https://dataverse.harvard.edu/dataset.xhtml?persistentId=doi:10.7910/DVN/DBW86T |
| FairVLMed | SLO fundus | 8254 | 1746 | https://ophai.hms.harvard.edu/datasets/harvard-fairvlmed10k |
| Camelyon17 | Pathology | 3680 | 920 | https://camelyon17.grand-challenge.org/Data/ |
Appendix C Implementation Details
C.1 Prompt for GPT Score
Traditional exact match-based metrics such as tokenized F1 and BLEU are sensitive to language patterns, making them ineffective for evaluating synonyms, especially for open-ended questions. Consequently, semantically correct answers generated by models not fine-tuned on the current dataset may receive disproportionately low scores in token similarity-based comparisons due to variations in language usage and conventions. To this end, we assess the correctness of predictions using LLMs, enabling semantic-level comparisons that are agnostic to language patterns.
For open-ended questions, to evaluate model performance with more advanced and flexible metrics, we adopt a method inspired by Li et al. [li2023llava], where a large language model is prompted to act as an automatic judge. Specifically, we design an open-form evaluation prompt, in which the LLM is asked to assess a candidate response against the reference answer in terms of helpfulness, relevance, accuracy, and level of detail. The model is then instructed to output both (i) a numerical score ranging from 1 to 100, reflecting the overall quality, and (ii) a free-form justification explaining the reasoning behind the score. This procedure yields a GPT-score tailored to our task setting. For closed-ended questions, we assign a score of 0 or 100 based on whether its yes/no answer is correct. We note that the GPT score is equivalent to accuracy for closed-form questions.
C.2 Prompt for VQA Dataset Construction from Clinical Notes
FairVLMed is an SLO fundus imaging dataset comprising disease classification labels and paired image–clinical note records. We leverage this dataset for both disease diagnosis and visual question answering (VQA). For the VQA evaluation, we employed GPT-4 to generate clinically oriented questions and reference answers for each image–clinical note pair. Specifically, for every pair we constructed one open-ended question and one closed-ended (yes/no) question, together with their corresponding answers. The prompting template used to elicit these question–answer pairs is provided below.
Appendix D Detailed Results
In this section, we present the detailed numerical results of this paper.
D.1 Diagnostic
| Model | Metric | Camelyon17 | CheXpert | GF3300 | HAM10000 | FairVLMed | PAPILA |
| BLIP2-2.7b | AUROC | 51.74% | 56.94% | 36.83% | 54.14% | 39.16% | 40.19% |
| CI | 47.97% | 46.16% | 31.83% | 50.14% | 36.43% | 17.20% | |
| CI | 55.59% | 68.20% | 41.25% | 57.88% | 41.77% | 61.07% | |
| AUC | 1.94% | 5.57% | 2.36% | 2.01% | 1.36% | 11.26% | |
| BioMedCLIP | AUROC | 77.17% | 60.71% | 47.86% | 56.79% | 54.35% | 76.12% |
| CI | 74.17% | 50.23% | 43.52% | 52.89% | 51.69% | 53.10% | |
| CI | 80.29% | 70.96% | 52.74% | 60.33% | 56.99% | 93.23% | |
| AUC | 1.56% | 5.47% | 2.41% | 1.88% | 1.37% | 10.53% | |
| CLIP | AUROC | 52.96% | 40.95% | 47.53% | 46.47% | 58.02% | 59.81% |
| CI | 49.27% | 31.68% | 42.85% | 42.76% | 55.23% | 37.87% | |
| CI | 56.83% | 50.67% | 52.21% | 50.51% | 60.81% | 80.79% | |
| AUC | 1.93% | 4.96% | 2.39% | 2.01% | 1.37% | 11.20% | |
| MedCLIP | AUROC | 76.39% | 90.60% | 38.15% | 48.66% | 52.14% | 36.41% |
| CI | 73.25% | 84.78% | 33.77% | 44.63% | 49.38% | 13.24% | |
| CI | 79.40% | 95.39% | 42.72% | 52.94% | 54.93% | 59.00% | |
| AUC | 1.58% | 2.75% | 2.27% | 2.09% | 1.42% | 11.67% | |
| MedSigLIP | AUROC | 42.14% | 62.29% | 54.55% | 52.70% | 59.66% | 56.97% |
| CI | 38.54% | 52.46% | 49.91% | 48.81% | 56.98% | 30.03% | |
| CI | 45.89% | 72.21% | 59.33% | 56.35% | 62.18% | 84.84% | |
| AUC | 1.93% | 4.96% | 2.41% | 1.91% | 1.38% | 14.11% | |
| PLIP | AUROC | 74.48% | 46.82% | 36.35% | 45.76% | 46.14% | 52.01% |
| CI | 71.29% | 36.50% | 32.05% | 41.89% | 43.38% | 32.60% | |
| CI | 77.35% | 57.85% | 40.84% | 49.58% | 48.78% | 72.93% | |
| AUC | 1.57% | 5.46% | 2.28% | 1.96% | 1.36% | 10.49% | |
| PubMedCLIP | AUROC | 33.04% | 76.44% | 41.34% | 41.25% | 46.02% | 37.59% |
| CI | 29.48% | 68.00% | 36.74% | 37.42% | 43.41% | 11.18% | |
| CI | 36.64% | 83.44% | 46.30% | 44.79% | 48.58% | 63.27% | |
| AUC | 1.82% | 3.84% | 2.37% | 1.90% | 1.34% | 12.95% | |
| SigLIP | AUROC | 59.81% | 53.42% | 55.23% | 46.99% | 53.02% | 51.77% |
| CI | 56.20% | 43.46% | 50.38% | 43.46% | 50.32% | 32.76% | |
| CI | 63.69% | 62.69% | 59.67% | 50.62% | 55.83% | 73.44% | |
| AUC | 1.84% | 4.98% | 2.33% | 1.86% | 1.39% | 10.43% |
| Model | Metric | Camelyon17 | CheXpert | GF3300 | HAM10000 | FairVLMed | PAPILA |
| BLIP2-2.7b | AUROC | 98.13% | 86.02% | 81.13% | 87.59% | 71.84% | 87.00% |
| CI | 97.44% | 79.75% | 77.49% | 85.87% | 69.35% | 72.51% | |
| CI | 98.74% | 91.43% | 84.49% | 89.41% | 74.32% | 97.62% | |
| AUC | 0.34% | 2.94% | 1.78% | 0.90% | 1.27% | 6.37% | |
| BioMedCLIP | AUROC | 98.57% | 88.52% | 81.72% | 86.54% | 71.87% | 72.34% |
| CI | 97.94% | 83.02% | 78.45% | 84.51% | 69.43% | 52.48% | |
| CI | 99.10% | 93.32% | 85.16% | 88.34% | 74.33% | 90.00% | |
| AUC | 0.30% | 2.60% | 1.72% | 1.02% | 1.25% | 9.58% | |
| CLIP | AUROC | 96.79% | 82.14% | 79.16% | 83.88% | 68.38% | 76.12% |
| CI | 95.72% | 75.04% | 75.49% | 81.76% | 65.78% | 60.64% | |
| CI | 97.71% | 88.86% | 82.55% | 85.93% | 70.87% | 89.80% | |
| AUC | 0.51% | 3.52% | 1.82% | 1.08% | 1.26% | 7.53% | |
| MedCLIP | AUROC | 81.70% | 86.94% | 70.98% | 73.54% | 61.09% | 81.56% |
| CI | 79.10% | 81.56% | 66.87% | 70.27% | 58.50% | 66.66% | |
| CI | 84.38% | 91.39% | 74.66% | 76.59% | 63.63% | 92.91% | |
| AUC | 1.39% | 2.44% | 2.05% | 1.64% | 1.38% | 7.01% | |
| MedSigLIP | AUROC | 98.60% | 88.91% | 76.53% | 86.20% | 73.53% | 87.94% |
| CI | 97.92% | 84.38% | 72.64% | 84.29% | 71.20% | 76.92% | |
| CI | 99.18% | 93.13% | 80.27% | 88.09% | 75.96% | 96.46% | |
| AUC | 0.32% | 2.29% | 1.95% | 0.98% | 1.23% | 5.12% | |
| PLIP | AUROC | 98.53% | 84.71% | 78.23% | 84.32% | 69.20% | 67.61% |
| CI | 97.77% | 78.32% | 74.48% | 82.25% | 66.64% | 49.80% | |
| CI | 99.16% | 90.59% | 81.81% | 86.22% | 71.77% | 84.45% | |
| AUC | 0.36% | 3.10% | 1.84% | 1.00% | 1.27% | 9.11% | |
| PubMedCLIP | AUROC | 95.59% | 84.34% | 77.15% | 84.11% | 69.10% | 51.77% |
| CI | 94.34% | 76.84% | 73.59% | 81.83% | 66.50% | 30.71% | |
| CI | 96.69% | 90.71% | 80.63% | 86.26% | 71.64% | 73.44% | |
| AUC | 0.60% | 3.55% | 1.85% | 1.11% | 1.28% | 10.77% | |
| SigLIP | AUROC | 97.21% | 85.37% | 77.20% | 87.34% | 71.45% | 71.16% |
| CI | 96.15% | 78.43% | 73.53% | 85.61% | 68.97% | 54.90% | |
| CI | 98.08% | 90.64% | 80.70% | 89.15% | 74.08% | 85.43% | |
| AUC | 0.48% | 3.14% | 1.85% | 0.91% | 1.25% | 7.67% |
| Model | Metric | Camelyon17 | CheXpert | GF3300 | HAM10000 | FairVLMed | PAPILA |
| BLIP2-2.7b | AUROC | 98.25% | 85.86% | 81.39% | 87.53% | 71.99% | 94.09% |
| CI | 97.57% | 79.44% | 77.85% | 85.82% | 69.47% | 86.01% | |
| CI | 98.83% | 91.39% | 84.72% | 89.38% | 74.49% | 99.06% | |
| AUC | 0.32% | 2.97% | 1.77% | 0.91% | 1.27% | 3.33% | |
| BioMedCLIP | AUROC | 98.70% | 88.37% | 81.67% | 86.99% | 71.92% | 68.32% |
| CI | 98.13% | 82.94% | 78.43% | 85.02% | 69.47% | 47.63% | |
| CI | 99.20% | 93.12% | 85.00% | 88.81% | 74.36% | 87.48% | |
| AUC | 0.28% | 2.61% | 1.70% | 0.99% | 1.26% | 10.27% | |
| CLIP | AUROC | 96.80% | 82.20% | 79.23% | 83.76% | 68.62% | 80.14% |
| CI | 95.80% | 74.84% | 75.52% | 81.73% | 66.10% | 65.25% | |
| CI | 97.70% | 88.94% | 82.61% | 85.77% | 71.07% | 92.55% | |
| AUC | 0.49% | 3.54% | 1.80% | 1.10% | 1.27% | 7.27% | |
| MedCLIP | AUROC | 81.84% | 86.69% | 70.67% | 74.11% | 61.59% | 55.32% |
| CI | 79.14% | 81.55% | 66.53% | 70.78% | 58.94% | 31.31% | |
| CI | 84.50% | 91.14% | 74.60% | 76.97% | 64.16% | 81.34% | |
| AUC | 1.38% | 2.42% | 2.09% | 1.59% | 1.37% | 12.85% | |
| MedSigLIP | AUROC | 98.59% | 88.80% | 76.27% | 86.36% | 73.57% | 91.25% |
| CI | 97.92% | 84.18% | 72.43% | 84.49% | 71.23% | 80.00% | |
| CI | 99.17% | 93.00% | 80.08% | 88.24% | 76.03% | 98.58% | |
| AUC | 0.33% | 2.30% | 1.96% | 0.97% | 1.22% | 4.63% | |
| PLIP | AUROC | 98.58% | 84.76% | 78.19% | 84.29% | 69.03% | 71.39% |
| CI | 97.86% | 78.23% | 74.55% | 82.29% | 66.47% | 57.31% | |
| CI | 99.20% | 90.56% | 81.74% | 86.20% | 71.66% | 85.10% | |
| AUC | 0.35% | 3.12% | 1.84% | 1.02% | 1.27% | 7.27% | |
| PubMedCLIP | AUROC | 95.56% | 84.51% | 76.80% | 83.97% | 69.41% | 79.43% |
| CI | 94.37% | 76.98% | 73.10% | 81.72% | 66.81% | 65.89% | |
| CI | 96.65% | 90.90% | 80.45% | 86.12% | 71.95% | 91.52% | |
| AUC | 0.59% | 3.51% | 1.87% | 1.12% | 1.28% | 6.60% | |
| SigLIP | AUROC | 97.20% | 85.30% | 77.74% | 86.69% | 71.06% | 71.39% |
| CI | 96.13% | 78.45% | 73.99% | 84.88% | 68.61% | 53.35% | |
| CI | 98.07% | 90.61% | 81.16% | 88.54% | 73.65% | 87.68% | |
| AUC | 0.48% | 3.16% | 1.84% | 0.94% | 1.25% | 8.62% |
D.2 VQA
Off-the-shelf Fine-tuned (ML) Metric Lingshu Qwen25-VL Qwen2-VL Lingshu Qwen25-VL Qwen2-VL Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric MedGemma Gemma3 MedGemma Gemma3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric LLaVA-Med LLaVA-1.5 LLaVA-Med LLaVA-1.5 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric VILA-M3 VILA VILA-M3 VILA Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric InternVL3 InternVL3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric Lingshu Qwen25-VL Qwen2-VL Lingshu Qwen25-VL Qwen2-VL Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric MedGemma Gemma3 MedGemma Gemma3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric LLaVA-Med LLaVA-1.5 LLaVA-Med LLaVA-1.5 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric VILA-M3 VILA VILA-M3 VILA Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric InternVL3 InternVL3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric Lingshu Qwen25-VL Qwen2-VL Lingshu Qwen25-VL Qwen2-VL Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric MedGemma Gemma3 MedGemma Gemma3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric LLaVA-Med LLaVA-1.5 LLaVA-Med LLaVA-1.5 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric VILA-M3 VILA VILA-M3 VILA Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric InternVL3 InternVL3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric Lingshu Qwen25-VL Qwen2-VL Lingshu Qwen25-VL Qwen2-VL Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric MedGemma Gemma3 MedGemma Gemma3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric LLaVA-Med LLaVA-1.5 LLaVA-Med LLaVA-1.5 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric VILA-M3 VILA VILA-M3 VILA Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
Off-the-shelf Fine-tuned (ML) Metric InternVL3 InternVL3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score